OpenStudy (anonymous):

Decide whether the function is an exponential growth or exponential decay function and find the constant percentage rate of growth or decay. f(x) = 2063 ⋅ 0.9953x Exponential decay function; -0.47% Exponential decay function; 0.0047% Exponential growth function; 0.0047% Exponential growth function; -0.47%

OpenStudy (anonymous):

*0.9953^x

OpenStudy (anonymous):

decay factor = 0.9953 rate of decay = (1 - 0.9953)*100 = -

OpenStudy (anonymous):

it is decay because you have 0.9953 if it is 1.02 it is growth

OpenStudy (anonymous):

okay so it is either a or b now how do i figure out which one it is?

OpenStudy (anonymous):

why B ? (1-0.9953)*100=0.47

OpenStudy (anonymous):

btw for decay the percentage is always negative